OpenStudy (anonymous):

f(x)=7x^3-3x^2+3 using f(a+h)-f(a)/h

OpenStudy (anonymous):

I started with 7(a+h)^3-3(a+h)^2+3/h But from there I do not know how to FOIL out (a+h)^3 properly. I would normally foil it out as if it were squared and then multiply the last (a+h) to what I got when it was expanded from squared form

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

f(x)=7x^3-3x^2+3 f(a+h)=7(a+h)^3-3(a+h)^2+3 f(a+h)=7(a+h)(a+h)^2-3(a+h)^2+3 f(a+h)=7(a+h)(a^2+2ah+h^2)-3(a^2+2ah+h^2)+3 f(a+h)=7[a(a^2+2ah+h^2)+h(a^2+2ah+h^2)]-3(a^2+2ah+h^2)+3 f(a+h)=7[a^3+2a^2h+ah^2+a^2h+2ah^2+h^3]-3(a^2+2ah+h^2)+3 f(a+h)=7a^3+14a^2h+7ah^2+7a^2h+14ah^2+7h^3-3a^2-6ah-3h^2+3 f(a+h)=7a^3+14a^2h+7ah^2+7a^2h+14ah^2+7h^3-3a^2-6ah-3h^2+3 f(a+h) = 7a^3+21a^2h+21ah^2+7h^3-3a^2-6ah-3h^2+3

OpenStudy (anonymous):

from there I factor out an ah from the numerator which then the h i pulled out cancels with the h from the denominator and I am left with a(7a^2+21a-21h+7h^2-3a-3h-3

OpenStudy (anonymous):

*+21a+21h*

jimthompson5910 (jim_thompson5910):

well, f(a+h) - f(a) turns into (7a^3+21a^2h+21ah^2+7h^3-3a^2-6ah-3h^2+3) - (7a^3-3a^2+3) 7a^3+21a^2h+21ah^2+7h^3-3a^2-6ah-3h^2+3 - 7a^3+3a^2-3 21a^2h+21ah^2+7h^3-6ah-3h^2 then you factor out h to get h(21a^2+21ah+7h^2-6a-3h)

jimthompson5910 (jim_thompson5910):

this is all in the numerator of course

OpenStudy (anonymous):

- (7a^3-3a^2+3) where does this part come from?

jimthompson5910 (jim_thompson5910):

that's -f(a) well after you replace f(a) with 7a^3-3a^2+3

OpenStudy (anonymous):

ohhhhh I was leaving that off...hang on a moment let me back up a little bit

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

@jimthompson_5910 > so the answer over all is: 21a^2+21ah+7h^2-6a-3h???

jimthompson5910 (jim_thompson5910):

that is correct

OpenStudy (anonymous):

wonderful! Thank you

jimthompson5910 (jim_thompson5910):

yw